OpenStudy (anonymous):

Solve for x: 3|2x-2|+6 = 18

OpenStudy (aaronandyson):

do the bracket times 3

OpenStudy (aaronandyson):

so its 6x - 6 + 6 = 18

OpenStudy (aaronandyson):

-6 +6 = 0

OpenStudy (aaronandyson):

6x = 18

OpenStudy (aaronandyson):

x = 18/6

OpenStudy (aaronandyson):

which number time 6 gives 18 is the answer

OpenStudy (anonymous):

x=3

OpenStudy (aaronandyson):

its 3

OpenStudy (aaronandyson):

hence x = 3

OpenStudy (aaronandyson):

NO WAIT I MADE A MISTAKE

OpenStudy (anonymous):

Here's my question options: x=3, x=-3 x=3, x=-1 x=1, x=-1 x=-3, x=1

OpenStudy (aaronandyson):

wait

OpenStudy (aaronandyson):

that -2 become +2 as in the mode nothing can be negative

OpenStudy (aaronandyson):

so thats 6x + 6 + 6 =18 6x + 12 = 18 6x = 6 x = 1

OpenStudy (aaronandyson):

so its D

OpenStudy (anonymous):

Okay, I think I get it now. Thank you!